OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (–2, 6) and has a slope of –1?

OpenStudy (bahrom7893):

y-5=-(x+2) y+x=3

OpenStudy (bahrom7893):

I think that's standard form

OpenStudy (anonymous):

True but you are wrong. Maybe it is y+x=4?

OpenStudy (bahrom7893):

y=-x-2+5 y+x=5-2 y+x=3

OpenStudy (anonymous):

x + y = –4 x + y = 4 x – y = –4 x – y = 4

OpenStudy (anonymous):

my only options

OpenStudy (bahrom7893):

y-6

OpenStudy (bahrom7893):

darn it..

OpenStudy (bahrom7893):

it's y+x=4.. sorry for some reason i wrote 5 instead of 6

OpenStudy (anonymous):

its okay hah thanks!